OpenStudy (anonymous):

Find the average rate of change from x = 10 to x = 17 for the function f(x) = 0.01(2)^x and select the correct answer below. A)10.24 B)185.78 C)1300.48 D)1310.72

OpenStudy (anonymous):

@jim_thompson5910 ???

jimthompson5910 (jim_thompson5910):

if x = 10, then f(x) = ???

OpenStudy (anonymous):

17?

jimthompson5910 (jim_thompson5910):

replace x with 10 and evaluate the function

jimthompson5910 (jim_thompson5910):

f(x) = 0.01(2)^x f(10) = 0.01(2)^10 f(10) = ???

OpenStudy (anonymous):

10.24?

OpenStudy (anonymous):

Just a guess, but A 10.24

jimthompson5910 (jim_thompson5910):

10.24 isn't the answer, but it helps get us there

OpenStudy (anonymous):

I am not really sure though

OpenStudy (anonymous):

:(

jimthompson5910 (jim_thompson5910):

if x = 17, then f(x) = ???

OpenStudy (anonymous):

0.01(2)

OpenStudy (anonymous):

? @jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

f(x) = 0.01(2)^x f(17) = 0.01(2)^17 f(17) = ???

OpenStudy (anonymous):

sorry, I don't know

OpenStudy (anonymous):

13?

jimthompson5910 (jim_thompson5910):

what is 0.01(2)^17 equal to

OpenStudy (anonymous):

WAIT its the answer D?

jimthompson5910 (jim_thompson5910):

nope, it helps us get there though

jimthompson5910 (jim_thompson5910):

f(10) = 10.24 f(17) = 1310.72

jimthompson5910 (jim_thompson5910):

now evaluate \[\Large \frac{f(17)-f(10)}{17-10}\]

OpenStudy (anonymous):

-17 f(10) +f(17) -10?

OpenStudy (anonymous):

its right?

jimthompson5910 (jim_thompson5910):

\[\Large \frac{f(17)-f(10)}{17-10} = \frac{1310.72-10.24}{17-10} = ???\]

OpenStudy (anonymous):

1300

OpenStudy (anonymous):

?

OpenStudy (anonymous):

1300.1*

jimthompson5910 (jim_thompson5910):

what is 1310.72-10.24 equal to? what is 17-10 equal to? divide those two results

OpenStudy (anonymous):

what is 1310.72-10.24 equal to 1300.48 what is 17-10 equal to 7

OpenStudy (anonymous):

im right?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

now divide 1300.48/7

OpenStudy (anonymous):

185.7?

jimthompson5910 (jim_thompson5910):

185.78

OpenStudy (anonymous):

185.78

OpenStudy (anonymous):

yea sorry

jimthompson5910 (jim_thompson5910):

yeah

jimthompson5910 (jim_thompson5910):

its ok

OpenStudy (anonymous):

soits b?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thank you, can you check something if make sure if im right?

OpenStudy (anonymous):

Identify the type of sequence shown in the table below and select the appropriate response. n f(n) 1 12 2 −36 3 108 4 −324 5 972

OpenStudy (anonymous):

A)Arithmetic sequence; common difference is −48 B)Arithmetic sequence; common difference is 144 C)Geometric sequence; common ratio is −3 D)Geometric sequence; common ratio is 3

OpenStudy (anonymous):

I THINK ITS C, but im not sure

OpenStudy (anonymous):

im right?

jimthompson5910 (jim_thompson5910):

why C?

OpenStudy (anonymous):

i dont know somebody give the answer

OpenStudy (anonymous):

@jim_thompson5910 ???

OpenStudy (anonymous):

are u there? @jim_thompson5910 ??/

jimthompson5910 (jim_thompson5910):

have a look at this article https://purplemath.com/modules/series3.htm